### v6

# Towards interactive evaluations for interaction harms in human-AI systems

Abstract
Current AI evaluation paradigms that rely on static, model-only tests fail to capture harms that emerge through sustained human-AI interaction. As interactive AI systems, such as AI companions, proliferate in daily life, this mismatch between evaluation methods and real-world use becomes increasingly consequential. We argue for a paradigm shift toward evaluation centered on interactional ethics, which addresses risks like inappropriate human-AI relationships, social manipulation, and cognitive overreliance that develop through repeated interaction rather than single outputs. Drawing on human-computer interaction, natural language processing, and the social sciences, we propose principles for evaluating generative models through interaction scenarios and human impact metrics. We conclude by examining implementation challenges and open research questions for researchers, practitioners, and regulators integrating these approaches into AI governance frameworks.
1 Introduction
Artificial intelligence (AI) model evaluations, broadly defined as systematic empirical assessments of AI models’ capabilities and behaviors, have become central to developers’ and regulators’ efforts to ensure that AI systems are sufficiently safe. Governments around the world have emphasized the importance of conducting model evaluations for various risks from discrimination to cybersecurity risks (UK Government, 2024; The White House, 2023); AI labs, including OpenAI with its Preparedness Framework and Anthropic with its Responsible Scaling Policy (OpenAI, 2023; Anthropic, 2023), propose using model evaluations to monitor and mitigate misuse and catastrophic risks; and academic researchers are developing evaluation datasets at unprecedented rates (Röttger et al., 2024). This positions model evaluations as integral to a range of important decisions on the safe development and deployment of AI systems.
The growing importance of model evaluations has been accompanied by increased scrutiny, with researchers both the unique challenges of evaluating generative AI compared to traditional machine learning systems, and broader concerns about validity, replicability, and quality (Weidinger et al., 2023; Liao and Xiao, 2023; Raji et al., 2022; Wallach et al., 2025). A prominent thread in these concerns is the disconnect between current evaluation approaches and real-world use of AI systems — mostly in the form of large language models (LLMs) — today. Despite some notable exceptions, the majority of current evaluations primarily rely on static, isolated tests that assess models based on their responses to individual prompts (Chang et al., 2024; Weidinger et al., 2023). Yet, this approach is at odds with real-world AI use, where applications increasingly depend on sustained back-and-forth interaction — from AI romantic companions engaging millions of users in dynamic conversations to agentic AI systems that take actions on users’ behalf.
This mismatch between evaluation methods and real-world use has become increasingly consequential as interactive AI systems proliferate homes, schools, and workplaces. Many of policymakers’, developers’, and the public’s most pressing concerns about AI systems — such as inappropriate human-AI relationships, social influence and manipulation, and overreliance on cognitive tasks — all arise from patterns of repeated interaction over time, not from single outputs. Consider a college student using an AI system as a companion for daily conversations over several months. While each individual response from the AI passes standard safety evaluations, the system could gradually shapes the student’s decision-making and emotional dependencies in concerning way — a pattern of harm entirely missed by conventional evaluation methods. Although current evaluations can detect content risks, such as toxicity and bias in individual output,they miss more subtle risks that emerge through sustained contextual interaction (Alberts et al., 2024).
In this paper, we argue that current AI evaluation paradigms we argue that current AI evaluation paradigms, while valuable for many purposes, are unable to fully capture harms that emerge through extended human interaction with AI systems. Specifically, we call for a paradigm shift to assessing AI systems through the lens of interactional ethics, focusing on interaction harms that arise through sustained engagement with AI systems as social and relational actors (Alberts et al., 2024).  [^1] Our examples focus on LLMs and text interactions because they provide the richest available data for studying sustained human-AI engagement patterns. However, the principles discussed could be applied to multimodal systems as they mature. Human participation is already playing a growing role in AI research, with studies involving human data making up approximately 9% of papers at top computer science venues AAAI and NeurIPS from 2021 to 2023 (McKee, 2024). Concurrently, social science research is increasingly conducting large-scale experiments with AI systems to understand their impact on human behavior (Costello et al., 2024). Our approach integrates these complementary insights from human-computer interaction (HCI), natural language processing, and the social sciences.
We begin by analyzing how current evaluation approaches fall short in assessing interaction harms due to their (1) static nature, (2) assumption of a universal user experience, and (3) limited construct validity. Drawing on decades of HCI research, we then propose organizing principles for evaluating generative models, structured around designing interaction scenarios, measuring human impact, and determining appropriate forms of human participation. We conclude by outlining key challenges and opportunities for researchers, companies, and regulators seeking to implement interactive evaluations of generative AI systems, including considerations of scalability, standardization, and practical integration with existing frameworks.
2 An overview of the generative AI evaluation landscape
We begin by examining contemporary approaches to ethics and safety evaluations of generative AI systems — their methodologies, primary focus areas, and limitations.
2.1 The current state of AI safety evaluations
We follow existing work in adopting a wide definition for “safety” which encompasses model behaviors associated with various taxonomized harms (Weidinger et al., 2022; Shelby et al., 2023). Examples of these include different types of biases (Parrish et al., 2021), toxicity (Hartvigsen et al., 2022), and “dangerous capabilities” like persuasion and cybersecurity risks (Phuong et al., 2024; Li et al., 2024). Safety evaluations of generative AI systems build on a rich history of ethical considerations in the field of natural language processing (NLP), prior to the advent of large pre-trained models, where researchers have long grappled with issues of harm from language technologies (Dev et al., 2021).
Recent reviews of existing safety evaluations show that the majority are focused on evaluating individual model responses to prompts curated by researchers targeting various model behaviors (Röttger et al., 2024; Weidinger et al., 2023). When safety evaluations do involve human subjects or evaluate over multiple dialogue turns, they often take the form of “red teaming campaigns” or automated adversarial testing that assumes malicious user intent (Zhang et al., 2024; Perez et al., 2022). The reviews of the landscape highlight two main gaps: (1) a methodological gap with the absence of evaluations over multiple dialogue turns (i.e., multi-turn evaluations), and as a result, (2) a coverage gap since interaction harms such as many social and psychological harms require evaluations that go beyond assessing model behavior in isolated, single-turn interactions.
A small but growing body of work has begun addressing these gaps, by conducting studies that utilize production chat logs (Phang et al., 2025) or user simulations (Ibrahim et al., 2025; Zhou et al., 2024) to investigate risks related to affective use of AI systems. Our work builds on and extends these emerging approaches to interactive evaluation.
2.2 Critiques of current evaluations
With the growing adoption of model evaluations, research has increasingly pointed out validity issues in how they assess generative models. Some researchers have questioned the external validity of current evaluations, noting that benchmark tasks poorly mirror real-world use cases and may not capture how models actually behave outside controlled settings (Ouyang et al., 2023; Liao and Xiao, 2023). Others have shown that current evaluations lack sufficient construct validity, especially when operationalizing complex concepts like fairness (Raji et al., 2021; Blodgett et al., 2021). In broader methodological reflections, researchers have also argued that LLMs should not be evaluated using frameworks designed for assessing humans, since LLMs exhibit distinct sensitivities, for example to prompt variations (McCoy et al., 2023). These challenges have led researchers to advocate for an interdisciplinary approach that draws on diverse traditions: the social sciences’ emphasis on measurement validity (Wallach et al., 2025), HCI’s focus on bridging technical capabilities and social requirements (Liao and Xiao, 2023), and cognitive science’s frameworks for analyzing system objectives (McCoy et al., 2023).
2.3 Methods for studying human-computer interaction
HCI research methods offer diverse ways to involve humans in studying computational systems — from having them as research subjects who actively engage with systems to having them serve as external observers and annotators. Drawing on this insight, we distinguish between static and interactive approaches to model evaluation based on whether they capture dynamic adaptation across multiple dialogue turns and occasions (Lee et al., 2022). Static evaluations assess fixed inputs and outputs in isolation, regardless of whether humans or automated systems perform this assessment. In contrast, interactive evaluations capture multi-turn exchange sequences that reveal how model behavior evolves through conversation and adapts to user responses, and how users are impacted by these model behaviors. This distinction transcends who conducts the evaluation; automated tests with realistic user simulations can qualify as interactive if they capture these dynamic patterns and validate their impact on real users (Ibrahim et al., 2025), while human studies that examine only single-turn responses remain static in nature (Hackenburg et al., 2023). Interactive evaluations may be “controlled," where interactions occur in structured, lab-like settings to systematically study specific variables or “retrospective," where researchers analyze existing multi-turn interactions from production chat logs to identify patterns and correlations between model behaviors and reported user experiences (Hariton and Locascio, 2018; Kuniavsky, 2003).
3 Why current evaluations approaches are insufficient for assessing interaction harms
First, most current evaluations are static, measuring model behavior with single-turn inputs and outputs. This fails to capture real-world use, where people often interact with models over multiple dialogue turns and multiple sessions.
Unlike some concerns about problematic model outputs — such as toxic language or factual errors — interaction harms manifest as gradual changes in human behavior, beliefs, or relationships that develop through repeated interactions, making them difficult to detect through single-turn evaluations. These harms are characterized by their compositional nature across conversation turns: while individual model responses may appear benign in isolation, their cumulative effect through ongoing dialogue can lead to concerning outcomes.
This temporal dimension operates through several mechanisms. First, subtle effects can compound over time. For instance, a single biased response might have limited impact, but repeated exposure to subtle biases across multiple interactions can shape a user’s decision-making patterns in high-stakes scenarios such as hiring. Second, psychological dynamics emerge through sustained relationships. Highly empathetic model responses that seem harmless in one exchange may, through sustained dialogue, lead users to form inappropriate emotional attachments or dependencies (Phang et al., 2025). Third, problematic behaviors may only emerge after multiple turns. Recent work demonstrates that certain safety-critical behaviors, such as models expressing anthropomorphic desires, sometimes only appear after a few conversation turns rather than in initial responses (Ibrahim et al., 2025).
This multifaceted temporal dimension distinguishes interaction harms from purely output-level concerns and demands longitudinal evaluation approaches that can capture these dynamic, cumulative effects.
Second, current evaluations collapse the diversity of user groups into a single, “universal” user. This overlooks how various demographic groups engage with models uniquely, and how models may tailor their responses to distinct user populations in complex ways. Current evaluation approaches rely heavily on standardized datasets where prompts are typically written by researchers or online crowdworkers. While using controlled test sets is a necessary starting point for systematic evaluation, this methodology implicitly assumes uniform user interactions. However, user groups which vary by demographics, domain expertise, technical knowledge, and psychological state (Liao and Sundar, 2022; Ibrahim et al., 2024) may exhibit different patterns of system interaction and use.
Beyond explicit customization features, models are developing increasingly sophisticated forms of implicit user modeling, forming internal representations that shape their responses (Chen et al., 2024; Staab et al., 2023). As a result, models sometimes engage in emergent forms of adaptation, dynamically adjusting their behavior based on perceived user attributes (Viégas and Wattenberg, 2023). While some adaptations may seem benign — like mirroring a user’s vocabulary as the interaction progresses — there is also evidence of concerning patterns. For instance, studies have found models varying their refusals of dangerous queries based on perceived user identity or persona (Ghandeharioun et al., 2025). Similarly, when users from different linguistic backgrounds employ various dialects of English or code-switch between languages, some empirical evidence shows that LLMs exhibit biased responses — displaying increased toxicity toward African American English compared to Standard American English (Hofmann et al., 2024). Factoring in this interaction variance becomes even more critical when evaluating generative models, which introduce new dimensions of uncertainty through their adaptive and probabilistic behaviors (Jin et al., 2024).
Third, current evaluations lack construct validity for interaction harms, as while they may measure AI capabilities, they cannot measure AI’s impact on human perception and behavior. A different set of metrics are needed to directly assess this impact.
Static evaluation methods can effectively identify harmful content like toxic or illegal content (Wallach et al., 2025). However, they fundamentally cannot address interaction harms where the harm lies not in the content itself but in its effect on users. Unlike identifying illegal content, where the evaluation goal is complete once the content is classified according to established criteria, interaction harms require establishing causal links between specific linguistic patterns and measurable human outcomes. For example, simply identifying that a response contains manipulative language does not confirm whether it actually influences user decision-making.
Thus, to establish construct validity for interaction harms, quantifying specific human impacts such as shifts in user beliefs, decision-making patterns, affective states, and dependency levels is needed. Importantly, recent work has shown that once these relationships between model behaviors and human impacts are validated through human experiments, the identified patterns may be repurposed as efficient static tests with empirically verified links to real-world harms (Ibrahim et al., 2025).
4 Towards better evaluations of interaction harms
The growing adoption of AI systems in daily life demands evaluation methods that can capture the nuanced human-AI interaction dynamics and their potential harms. Drawing on established methodologies for studying HCI and addressing the limitations identified in the previous section, we propose three organizing principles for developing interactive evaluations. We structure these principles around key challenges: designing ecologically valid scenarios, measuring human impact, and determining appropriate human participation:
  1. 1.
    Scenario design: developing more ecologically valid contexts that reflect real-world interaction patterns and user objectives
  2. 2.
    Impact measurement: establishing rigorous approaches and metrics for measuring how model behaviors impact human behaviors
  3. 3.
    Participation strategy: determining appropriate forms of human involvement to balance experimental control and costs with authentic interaction
Principle 1: design interaction scenarios based on user objectives and interaction modes
To systematically evaluate human-AI interaction, we need to consider two key dimensions: why users engage with these systems (their objectives) and how they interact with them (their interaction modes). In HCI, user goals or objectives have been shown to shape how they engage with systems and thus influence the outcomes of these interactions (Subramonyam et al., 2024). Therefore, we provide a categorization of harmful scenarios based on user objectives and affected parties, as shown in Table 1. These scenarios capture key patterns of harm from current generative AI systems that researchers and practitioners have observed (Mitchell, 2024).
Table 1: Possible harmful use scenarios and examples of each
Scenario Misuse Unintended harm: personal impact Unintended harm: external impact
Objective User intentionally uses model to inflict harm on another person or group of people User uses model, gets harmed in the process User uses model, unintentionally harms another person or group of people
Affected parties External subjects User External subjects
Example(s) User utilizes model capabilities for planning a biological attack User overrelies on empathetic model for mental health support, delaying professional help User trusts biased model judgment, making a discriminatory hiring decision
Beyond their objectives, users’ different patterns of engagement with AI systems shape how their interactions unfold. A single generative model can support multiple interaction modes, from simple question-answering to complex collaborative tasks. Drawing on observed use cases and literature reviews, we identify five prototypical modes of human-model interaction (Gao et al., 2024; Ouyang et al., 2023; Zhao et al., 2024; Händler, 2023; Collins et al., 2023), visualized in Figure 1:
  • Collaborative: human and model work in tandem towards completing joint goal-oriented tasks (e.g., human and model iteratively write and refine a report together)
  • Directive: human delegates specific tasks to the model for independent completion (e.g., human instructs model to generate a marketing campaign)
  • Assistive: model provides supporting input while human maintains primary agency (e.g., model suggests edits while human writes a document)
  • Cooperative: human and model make distinct contributions toward a shared goal without directly working together (e.g., model generates data visualizations while human writes analysis)
  • Explorative: human engages in open-ended interaction without specific task goals (e.g., casual conversation or creative brainstorming)
Figure 1: Taxonomy of human-AI interaction modes
Principle 2: identify the causal link between model behavior and human impact
To evaluate interaction harms in a valid way, we must trace how model behaviors influence human users. In Figure 2, we present an example of such a trace: underlying model properties shape observable model behaviors, which affect users through various pathways — from shaping moment-to-moment interactions to influencing deeper patterns of perception, decision-making, and behavior. Specifying the hypothesized pathways helps determine both what to manipulate in evaluations (model behaviors) and what to measure (human impact). We identify three key measurement targets:
  • Psychological impact: changes in user perceptions, attitudes and beliefs, and affective states
  • Behavioral impact: changes in user actions during and after interaction
  • Interaction outputs: quality and characteristics of the produced interaction artefact
Figure 2: Example of a causal trace showing how model properties may influence human behavior and human-AI interaction outcomes. Such traces help identify key measurement points for evaluation.
These measurement targets can be assessed through both self-reported measures — including validated scales for attitudes, perceptions, and affective states — and behavioral measures that track concrete actions like response time, task accuracy, and engagement patterns (Damacharla et al., 2018; Coronado et al., 2022; Gordon et al., 2021). In Table 2, we present a non-exhaustive set of these metrics for assessing both human impact and human-AI performance.
Table 2: Metrics to evaluate human-AI interactions
Evaluation target Description Metrics
Psychological impact User affective states, or perceptions and attitudes toward the model or the interaction Usability metrics, user satisfaction surveys, psychometrically validated measures for a specific construct or harm (e.g., Decision Regret Scale (Brehaut et al., 2003))
Behavioral impact (measured during the interaction) Observable patterns of behavior recorded during the interaction Number of queries users made, number of revisions users made, time between queries
Behavioral impact (measured after interaction) Observable behaviors assessed following the interaction Adherence to AI suggestions (e.g., choosing to donate when the AI recommends it), disclosing personal information when asked by the AI
Interaction output Objective quality of output, assessed either automatically (e.g., performance accuracy) or using third-party evaluators (human or LLM) Consistency of a summary given a document, success rate for completing a task
Principle 3: structure human participation to balance validity and practicality
Evaluating human-AI interaction can involve varying degrees of human participation: from controlled studies with live participants, to analyses of chat logs between users and deployed systems, to automated user simulations. Each approach offers distinct trade-offs in terms of ecological validity, experimental control, cost, and ethical considerations.
Chat logs from real-world interactions provide unparalleled ecological validity and scale. These naturally occurring conversations capture authentic user behaviors and emergent patterns that would be difficult to replicate in controlled settings (Tamkin et al., 2024). The massive volume of existing logs also enables analysis of rare events and edge cases that controlled studies might miss. Recent work has paired analyses of millions of chat logs of user conversations with ChatGPT with longitudinal surveys of those users to examine emotional well-being, showing a path forward for using chat logs while also assessing real user impact (Phang et al., 2025). However, these observational data is often proprietary, difficult to access, raises serious privacy concerns, and may limit our ability to test specific hypotheses or establish causal relationships (Reuel et al., 2024; Zhao et al., 2024; Ouyang et al., 2023).
Controlled human subject studies, while more resource-intensive, allow for the systematic examination of specific interaction patterns and outcomes. These studies are particularly valuable when we need to administer psychological measures or capture behavioral changes over time. However, their artificial nature and high costs make them better suited for targeted investigation of specific hypotheses rather than broad-scale evaluation.
Existing research on user simulations has utilized LLMs to simulate believable human behaviors using a range of architectures from demographic-based models to simulations informed by qualitative interviews (Park et al., 2024, 2023). Recent work has used such simulations to evaluate safety risks, utilizing human experiments to validate the evaluations (Ibrahim et al., 2025; Zhou et al., 2024). Such simulations can be particularly valuable for exploring scenarios that would be ethically challenging to test with real users. Unlike real human participants, who naturally shape conversations in diverse ways, simulations also allow for controlled trajectories, making it easier to analyze how interactions evolve across different scenarios. However, their limitations in believability, accuracy, and diversity require further study, as recent work has shown that they misportray and flatten representations of marginalized identity groups (Wang et al., 2025; Agnew et al., 2024).
5 Open challenges and ways forward for interactive evaluations
In this paper, we motivate the need to investigate human-AI interaction dynamics that current evaluations miss. However, implementing such evaluations at scale presents several concrete challenges. These range from ethical questions about studying potentially harmful interactions, to practical needs for research infrastructure, to methodological questions about producing actionable insights for stakeholders. While our design principles outline key considerations for scenarios, measurements, and participation strategies, advancing these methods requires addressing key open questions. Here, we identify specific challenges and opportunities where concentrated research effort could significantly advance our ability to evaluate increasingly interactive AI systems.
How can we ethically work with human participants on studying harms? When are user simulations appropriate replacements?
Studying interaction harms poses inherent ethical challenges: we need to understand potentially harmful dynamics without exposing participants to those same harms. This requires careful consideration of when and how to involve human participants. For participant protection, AI researchers should adopt existing practices from fields experienced in providing adequate participant training, thoughtful debriefing, and active feedback collection to mitigate experimental harms (McKee, 2024). Simulated human interactions offer a promising direction for studying scenarios where minimizing harm is challenging, such as young people’s interactions with AI systems or emotional attachment to AI systems. Recent work suggests these simulations can believably represent human behavior (Park et al., 2023, 2024), but more work is needed to understand how to best capture diverse user behaviors and develop ethical guidelines for deploying user simulations in safety evaluations (Agnew et al., 2024; Anthis et al., 2025).
How can we improve researcher access to data for understanding interaction harms?
Current public datasets of human-AI conversations, while valuable, fall short of meeting researchers’ needs. Large collections of chat logs like WildChat and ShareGPT offer millions of conversations but provide only a narrow window into AI usage, capturing self-selected interactions that may not represent typical user behaviors or patterns (Zhao et al., 2024; Ouyang et al., 2023). These datasets can be biased in terms of user demographics, use cases, and interaction styles, potentially leading to incomplete or incorrect conclusions about how people typically interact with AI systems. AI developers, such as large labs which provide commercial chatbot services, have access to the most comprehensive interaction data (Tamkin et al., 2024). While these labs have incentives to protect user privacy and proprietary information, developing privacy-preserving data sharing frameworks is necessary to enable researchers to access representative interaction data while maintaining user trust and consent.
What infrastructure do we need to facilitate interactive evaluations?
Many interactive evaluations share common stages: recruiting participants, collecting interaction data, and administering surveys and other tasks. Developing accessible protocols, guides, and standardized test suites — similar to those available for non-interactive evaluations — can support and facilitate an increase in interactive evaluations (UK AI Safety Institute, 2024; METR, 2023; Collins et al., 2023). Some existing platforms enable crowd-sourced interactive evaluations, but they face validity issues by gamifying safety testing (e.g., with tasks like "break the model in one minute") and reducing safety to vulnerability against jailbreaking (Angelopoulos et al., 2024). Instead, we need more platforms that standardize the foundational elements of human subject studies with AI while maintaining scientific rigor, reducing technical overhead for researchers, and enabling broader participation in AI evaluation.
How can interactive evaluations produce actionable findings that guide stakeholder decisions? What are the limitations of controlled studies in capturing broader impacts?
Randomized control trials (RCTs) have proven valuable for policy decisions by providing rigorous evidence, particularly in public health and social policy (Hariton and Locascio, 2018). Similarly, interactive evaluations aim to produce actionable findings about how model behaviors influence human perception and behavior. These insights can help stakeholders make better decisions about model deployment, safety mechanisms, and interaction design. However, we currently lack adequate visibility into how different stakeholders — from AI labs to government bodies — actually use evaluation results in their decision-making processes. Bridging this gap requires closer collaboration between researchers and decision-makers to ensure evaluation designs align with practical needs (Hardy et al., 2024).
Finally, we must be mindful of limitations; like RCTs, controlled evaluations of human-AI interaction may effectively measure individual-level effects (like individual manipulation or overreliance) while still missing broader systemic patterns. Thus, while measuring these individual impacts is crucial for understanding immediate safety concerns, they should be complemented with examinations of how AI systems reshape institutional structures, professional practices, and social arrangements — effects that emerge beyond individual interactions (Chater and Loewenstein, 2023).
6 Conclusion
As AI systems become increasingly integrated into our daily lives, moving beyond content risks to examine harms that emerge through sustained engagement is not just an academic exercise — it is essential for developing responsible technologies that prioritize human well-being. Several established industries, from medicine to automotives, have long recognized the need to study their technologies’ impact on users through extensive testing during development and after deployment (Wouters et al., 2020). As AI capabilities and applications continue to expand, we must similarly strengthen our investment in understanding these systems’ effects on human behavior and society at large.
Acknowledgments
We are incredibly grateful to Seth Lazar, Katy Glenn Bass, and our anonymous peer reviewers for helpful comments on the paper. We also thank the participants of the Knight First Amendment Institute workshop on AI and democratic freedoms for their feedback. Finally, thank you to Jamie Bernardi, Merlin Stein, Patrick Levermore, Kobi Hackenburg, Deep Ganguli, Ben Bucknall, Esin Durmus, and Christopher Summerfield for helpful comments on an earlier draft of the paper.